# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ 2.2 Controlled Direct Effect (CDE)
Table 1: A true partitioning exists for any $\mathbf{Z}$ (from [48]).
2.3 Mapping Causal Partitions to the Standard Fairness Model
To facilitate efficient structure learning, we leverage the causal partition taxonomy defined in [48] (Table 1). Given an exposure-outcome pair $\{X,Y\}$, any arbitrary variable set $\mathbf{Z}$ can be uniquely partitioned into exactly eight disjoint subsets (which may be empty) that are defined by the types of causal paths that they share with $X$ and $Y$ [^2] See [48] for formal definitions of these partitions and their proxies. Note that the notion of proxy variables is not relevant in this setting, as we are only concerned with true confounders and mediators. Narrowing the focus of structure learning to the relationships that are causally relevant to the exposure and outcome [12], this partition taxonomy provides convenient building blocks for local discovery. Mapping this partition taxonomy to the SFM, $\mathbf{Z}_{1}$ is a superset of confounders $\mathbf{C}$ and $\mathbf{Z}_{3}$ is a superset of mediators $\mathbf{M}$.
$X$$Y$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{4}$
Figure 2: The local subgraph identified by LD3. Here, members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ are parents of $Y$. Arrows   signify adjacency and   signify paths with arbitrarily many intervening variables. The SDC indicates whether the dotted edge from $X$ to $Y$ exists.
3 Local Discovery for Direct Discrimination (LD3)
Algorithm 1 provides a constraint-based discovery procedure for assessing the SDC and learning a valid adjustment set for weighted CDE estimation from observational data. The number of independence tests performed scales linearly with the number of nodes. Proofs of correctness are given in Appendix C.
Sufficient Conditions for Structure Learning
We assume causal Markov, faithfulness, and acyclicity. We do not impose parametric assumptions on causal functions nor distributional forms. As for all constraint-based methods, the independence test selected may impose its own parametric assumptions. When these are not well-justified, we recommend nonparametric tests (e.g., [28, 29, 99, 71]).
Asymptotic guarantees on partition label correctness, SDC correctness, and valid adjustment for unbiased weighted CDE estimation hold under Assumptions (A1) and (A2).
  1. (A1)
    $Y$ has no descendants in the observed variable set. This assumption is relatively straightforward to justify in settings where $Y$ is a terminal variable with respect to temporal ordering (e.g., when the outcome is death, or a decision made at a known time point).
  2. (A2)
    All parents of $Y$ are observed. Latent variables that are not parents of $Y$ are permissible.
Remark 3.1 (Notes on (A2)).
This assumption is a consequence of the sufficient conditions for identifiability of the CDE (Definition 2.5) [59, 62]. We exploit (A2) in our approach to evaluating the SDC [69], as it ensures that $X$ is conditionally $d$-separated from $Y$ when there is no direct path from $X$ to $Y$. It also ensures that no colliders in $\mathbf{Z}_{2\notin de(Y)}$ are retained for adjustment. If some parents of $Y$ are unobserved, the SDC might return a value of 1 and the weighted CDE might be non-zero even if there is no direct discrimination present.
We prove in Appendix C that latent variables not adjacent to $Y$ have no impact on correctness.
Theorem 3.2.
Latent variables that are not parents of $Y$ do not affect Algorithm 1.
Algorithm 1 LD3: Learning structural evidence of direct discrimination from observational data.
0:  Exposure $X$, outcome $Y$, variable set $\mathbf{Z}$, independence test, significance level $\alpha$.
0:  Adjustment set $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ and SDC results.
0:  (A1) $Y$ has no observed descendants and (A2) all parents of $Y$ are observed.
@@ 3 Local Discovery for Direct Discrimination (LD3)
3:     if $Z\perp\!\!\!\perp X\land Z\perp\!\!\!\perp Y$ then $Z\in\mathbf{Z}_{8}$
@@
6:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}^{\prime}\setminus\mathbf{Z}_{8}\cup\mathbf{Z}_{7}\cup\mathbf{Z}_{4}$
@@
9:  for $\forall\;Z_{4}\in\mathbf{Z}_{4}$ do
@@
11:  $\mathbf{A}_{\text{{DE}}}\leftarrow\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$
12:  if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ then $SDC\leftarrow 0$
@@
Given (A1), there are no descendants of $Y$ in $\mathcal{G}$. Thus, there are no $\mathbf{Z}_{6}$ nor $\mathbf{Z}_{2\in de(Y)}$. However, there may be $\mathbf{Z}_{2\notin de(Y)}$. Thus, $\mathcal{G}$ can contain the following seven (potentially empty) causal partitions with respect to $\{X,Y\}$: $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, $\mathbf{Z}_{4}$, $\mathbf{Z}_{5}$, $\mathbf{Z}_{7}$, and $\mathbf{Z}_{8}$. LD3 returns the subgraph pictured in Figure 2: all $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are directly adjacent to $Y$.
3.1 A Graphical Criterion for Identifying the Weighted CDE
Definition 3.3 (Graphical criterion for identifying the weighted CDE).
We define the set $\mathbf{A}_{\text{{DE}}}$, which contains all parents of the outcome. Under the causal partition taxonomy defined in [48],
@@ 3.1 A Graphical Criterion for Identifying the Weighted CDE
Theorem 3.4 ($\mathbf{A}_{\text{{DE}}}$ is a valid adjustment set for the weighted CDE).
The adjustment set $\mathbf{A}_{\text{{DE}}}$ satisfies the identification conditions for the weighted CDE (Definition 2.5) by blocking (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all frontdoor paths for $\{X,Y\}$.
Intuition.$\mathbf{A}_{\text{{DE}}}$ contains only and exactly all the parents of $Y$: confounders for $\{X,Y\}$ that are adjacent to $Y$ ($\mathbf{Z}_{1\in pa(Y)}$), mediators for $\{X,Y\}$ that are adjacent to $Y$ ($\mathbf{Z}_{3\in pa(Y)}$), and all parents of $Y$ that are marginally independent of $X$ ($\mathbf{Z}_{4\in pa(Y)}$). Thus, at least one member of every backdoor and frontdoor path is in $\mathbf{A}_{\text{{DE}}}$. This provides conditional $d$-separation of $X$ and $Y$ if and only if there is no edge $X\to Y$. Full proofs are provided in Appendix C.
Remark 3.5 (The role of $\mathbf{Z}_{4\in pa(Y)}$).
Note that $\mathbf{Z}_{3}$ and $Y$ can be confounded by $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, or $\mathbf{Z}_{4}$. Including $\mathbf{Z}_{4\in pa(Y)}$ in $\mathbf{A}_{\text{{DE}}}$ helps guarantee identifiability of the weighted CDE without requiring exact knowledge of confounding for $\mathbf{Z}_{3}$ and $Y$. As $\mathbf{Z}_{4}$ is not post-treatment, it is not forbidden from adjustment under the backdoor criterion [58]. It has been shown that adjusting for $\mathbf{Z}_{4}$ can reduce estimate variance in propensity score models [11], and conventional intuition dictates that adjustment should minimally constrain the exposure and maximally constrain the outcome [72]. Thus, there may be statistical benefits to including $\mathbf{Z}_{4\in pa(Y)}$, though we do not formally explore this here.
@@ 4 Related Works
Few works in causal discovery have centered on fairness objectives [9, 7]. To our knowledge, this work presents the first local causal discovery method that is specifically tailored for causal fairness analysis. Below, we summarize two lines of literature closely connected to this work.
Estimating Direct Causal Effects
Multiple estimands have been proposed for direct effect estimation, including the CDE [59], the NDE [59], and the counterfactual direct effect (Ctf-DE) [98]. The CDE is an interventional quantity, while the NDE and Ctf-DE are counterfactual quantities [62, 98]. The NDE is popular in causal fairness analysis [69] as it is guaranteed to admit a valid decomposition of the total causal effect into direct and indirect effects in the nonparametric setting [59]. However, when direct discrimination is the quantity of interest, the CDE is a valid measure [98]. Previously, the CDE has garnered more use in policy evaluation than the NDE [89]. In this work, we favor the weighted CDE as a direct discrimination metric in complex domains because it requires fewer untestable assumptions over the data generating process [82] and less prior structural knowledge than the NDE [59, 89]. This permits unbiased estimation in a wider range of structures, and a lower burden of causal discovery. We provide an extended justification in Appendix B.
Graphical Criteria for Direct Effect Identification
Sufficient and necessary conditions for unbiased identification vary by estimand, but generally require covariate adjustment [91]. Formal graphical rules can allow the valid adjustment set to be read directly from known causal structures, even in the presence of latent confounding [63, 49, 72]. Graphical criteria for valid adjustment have been proposed for diverse causal quantities, most famously the backdoor criterion for total effect estimation [58]. Often, multiple valid criteria exist per estimand. These can vary in the amount of prior structural knowledge that they require, and criteria that require less knowledge are easier to implement in practice [92, 30]. Shpitser & Pearl [81] present a complete graphical criterion for conditional interventional distributions. Pearl [59] defines a graphical criterion for identifying the NDE from observational data (Corollary 1). Discussion, debate, and confusion around the sufficient and necessary conditions for NDE identification have inspired a flurry of papers [37, 36], as summarized in Pearl [62]. Shpitser & VanderWeele [82] present a complete graphical criterion for the NDE that reconciles some debate, showing where certain common criteria coincide.
Graphical criteria for direct effect estimation are often challenging for practitioners to implement [82]. To increase the practicality of covariate adjustment in complex and low-knowledge domains, we advocate for the development of graphical criteria that are (1) relatively intuitive and practical to implement in applied research, and (2) are satisfied by knowledge that can be reliably learned through data-driven methods. Thus, we apply the causal partition taxonomy of [48] to provide local graphical intuition for fast and practical adjustment set discovery.
5 Empirical Validation
@@
We compare LD3 against global and local discovery baselines, with two evaluation methods per algorithm (one with post-processing adapted for this setting and one without, as described in Appendix D): PC (PC and PC) [38], [^3] https://github.com/keiichishima/pcalg LDP (LDP and LDPpa) [48], and LDECC (LDECC and LDECC) [31]. PC is a popular global discovery method that has spawned many extensions. LDP is asymptotically guaranteed to return a valid adjustment set for total effect estimation under the backdoor criterion [48]. LDECC returns the parents and children of a target variable [31]. All baselines are constraint-based to facilitate direct comparisons on the number of independence tests performed and evaluation with an oracle. Throughout, $n$ denotes data sample size (abbreviated with $k=1000$). All experiments were performed on a local MacBook Pro (M2 Pro Chip; 16 GB memory).
Figure 3: Mean runtime for the Asia and Sachs benchmarks from bnlearn ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. See Figure E.2 for total independence tests performed. Raw data are reported in Tables E.2 and E.4.
Adjustment Set Correctness
@@
As correct assessment of the SDC and unbiased estimation of the weighted CDE require that all parents of the outcome are correctly identified, we evaluated the ability of LD3 to recover the true parent set in randomly generated DAGs. Using an oracle independence test to simulate access to infinite data, we empirically validated LD3 on the adjacency matrics of random directed acyclic Erdős-Rényi graphs [23]. [^4] https://r.igraph.org/ LD3 perfectly recovered parent sets for randomly selected exposure-outcome pairs in 90 unique graphs. Ten graphs were generated per DAG size, with node counts in $[5,10,20,30,40,50,100,200,500]$. Parent F1, recall, and precision were 100%. Runtimes and total independence tests as node and edge cardinality scale are shown in Figure E.1.
Next, we assesed DAGs from the bnlearn Bayesian Network Repository [76]. Results for the Asia (Lung Cancer) dataset [45] are reported in Tables E.1 and E.2. Results for the Sachs protein expression dataset [74] are reported in Tables E.3 and E.4. For each benchmark, we selected one exposure-outcome pair for which SDC = 1 and SDC = 0. For Asia, LD3 and LDPpa had the highest F1 scores for both exposure-outcome pairs, with 100% F1 when SDC = 1 starting at $n=2.5k$ (Table E.1). For SDC = 0, LD3 improved with increasing sample size, while other methods decreased or peaked. [^5] See Table E.1 for elaboration. Notably, neither PC nor PC identified a single causal parent of the outcome. For Sachs, only LDECC attained 100% F1 for both pairs across all sample sizes, though LD3 and LDPpa were closely behind and still provided valid adjustment sets for the weighted CDE in all settings (Table E.3). The primary benefits of LD3 were efficiency and lax assumptions: the number of independence tests performed by LDECC and PC is worst-case exponential with respect to variable set size for random structures, LDP is worst-case quadratic, and LDP only guarantees correctness under the sufficient condition that at least one member of $\mathbf{Z}_{4}$ and $\mathbf{Z}_{5}$ are observed. Efficiency pitfalls for baselines are emphasized in Figure E.1 and Tables E.2 and E.4.
Weighted CDE Estimation
To evaluate the ability of LD3 to accurately assess the SDC and enable weighted CDE estimation, we tested linear-Gaussian instantiations of an 18-node DAG with an M-structure (i.e., a member of $\mathbf{Z}_{2\notin de(Y)}$) [22], a mediator chain, and a member of $\mathbf{Z}_{4}$ that confounds a mediator and $Y$ (Figure D.1). The weighted CDE was estimated using linear regression. $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall and SDC accuracy converged toward 100% as sample size increased, with the weighted CDE converging toward the true direct effect with low variance (Figures E.3, E.4).
6 Real-World Case Study: Fairness in Liver Transplant Allocation
We apply our proposed algorithm to a real-world case study in healthcare and compare its findings with past studies to illustrate the value that LD3 can bring in practice.
Allocation Under Scarcity

### v3

@@ 2.3 Mapping Causal Partitions to the SFM
Table 1: Adapted from Maasch et al. (2024).
3 Local Discovery for Direct Discrimination
In CFA, we can translate the fairness query “Is direct discrimination present?” into the graphical query “Is the protected attribute a parent of the outcome?” To answer this graphical query, we focus on two indicators of parentage: the SDC and WCDE, where the former can be directly answered by LD3 and the latter can be estimated by using the VAS returned by LD3.
The input to LD3 (Algorithm 1) is a variable set $\mathbf{Z}$ of unknown causal relation to the protected attribute $X$ and outcome $Y$. Instead of learning the causal graph, LD3 learns causal partition labels (Table 1). This local learning approach abstracts away structural information that is impertinent to direct discrimination detection, resulting in a computationally efficient discovery procedure. LD3 performs sequential CI tests to iteratively discover the partition label ($\widehat{\mathbf{Z}}$) of each variable in $\mathbf{Z}$. Leveraging these labels, LD3 uses a test of d-separation to evaluate SDC (Lines 12–13). By conditioning only on $\widehat{\mathbf{Z}}_{1,3\in pa(Y)}$, this CI test offers sample efficiency benefits relative to conditioning on all $\mathbf{Z}$. Additionally, LD3 returns a VAS for WCDE containing $pa(Y)\setminus X$.
Time Complexity
Constraint-based discovery methods are typically analyzed by the number of CI tests performed (Spirtes, Glymour, and Scheines 2001). For-loops at Lines 2–5, 7–8, and 9–10 of Algorithm 1 perform $O(|\mathbf{Z}|)$ tests each. All remaining lines perform a constant number of operations. Thus, the total number of CI tests is of $O(|\mathbf{Z}|)$, ensuring scalability in real-world CFA.
Algorithm 1 LD3
0:  Exposure $X$, outcome $Y$, variable set $\mathbf{Z}$, CI test of choice, significance level $\alpha$.
0:  Adjustment set $\mathbf{A}_{\text{{DE}}}$, SDC results.
0:  Sufficient conditions A1 and A2.
@@ 3 Local Discovery for Direct Discrimination
3:     if $Z\perp\!\!\!\perp X\land Z\perp\!\!\!\perp Y$ then $Z\in\widehat{\mathbf{Z}}_{8}$
@@
6:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}^{\prime}\setminus\widehat{\mathbf{Z}}_{8}\cup\widehat{\mathbf{Z}}_{5,7}\cup\widehat{\mathbf{Z}}_{4}$
@@
9:  for $\forall\;\widehat{Z}_{4}\in\widehat{\mathbf{Z}}_{4}$ do
@@
11:  $\mathbf{A}_{\text{{DE}}}\leftarrow\widehat{\mathbf{Z}}_{1\in pa(Y)}\cup\widehat{\mathbf{Z}}_{3\in pa(Y)}\cup\widehat{\mathbf{Z}}_{4\in pa(Y)}$
12:  if $X\perp\!\!\!\perp Y|\widehat{\mathbf{Z}}_{1\in pa(Y)}\cup\widehat{\mathbf{Z}}_{3\in pa(Y)}$ then $SDC\leftarrow 0$
@@
Sufficient Conditions for Structure Learning
We assume causal Markov, faithfulness, and acyclicity. We do not impose parametric assumptions on causal functions nor distributional forms. As for all constraint-based methods, the independence test selected may impose its own parametric assumptions. When these are not well-justified, we recommend nonparametric tests (e.g., Gretton et al. 2005, 2007).
Theorem 1.
Asymptotic guarantees on partitioning and SDC correctness hold under Assumptions A1 and A2. Given WCDE identifiability by Equation 5 (Remark 1), A1 and A2 are also sufficient for VAS discovery.
  1. A1
    $Y$ has no descendants in the observed variable set. This is satisfied when $Y$ is a terminal variable in the temporal ordering (e.g., when outcome is death, a policy or algorithmic decision made at a known time point, etc.).
  2. A2
    All parents of $Y$ are observed. Latent variables that are not parents of $Y$ are permissible. Thus, this is a milder condition than assuming causal sufficiency.
Proof of Theorem 1 is in Appendix B. Note that assumptions A1 and A2 are sufficient but not necessary. A1 has been previously used to facilitate parent and ancestor learning (Soleymani et al. 2022; Cai et al. 2023). LD3 learns causal partitions directly from data, without assumptions on temporal ordering other than A1 and $Y$ cannot cause $X$. Assumption A2 is a consequence of the sufficient conditions for CDE identifiability, as Definition 5 requires blocking all spurious and indirect paths into $Y$. A2 allows for unobserved variables that are not in $pa(Y)$, a more relaxed assumption than the causal sufficiency typically required in discovery (e.g., Spirtes, Glymour, and Scheines 2001; Zheng et al. 2018; etc.). We exploit A2 to evaluate the SDC, as it ensures that $X$ is conditionally $d$-separable from $Y$ when there is no direct path from $X$ to $Y$. It also ensures colliders ($\mathbf{Z}_{2\notin de(Y)}$) are removed from the adjustment set.
We empirically demonstrate failure modes and robustness to violations of A2 in Appendix D.1, showing that A2 is sufficient but not necessary. We prove in Appendix B that latent variables not adjacent to $Y$ do not impact correctness.
Theorem 2.
Latent variables that are not parents of $Y$ do not affect Algorithm 1.
$X$$Y$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{2}$$\mathbf{Z}_{4}$$\mathbf{Z}_{5}$$\mathbf{Z}_{7}$$\mathbf{Z}_{8}$?........(A)
$X$$Y$$\mathbf{Z}_{2}$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{4}$?(B)
Figure 2: LD3 assesses whether the edge $X\to Y$ exists. (A) Allowable partitions under A1 and A2. (B) Parents of $Y$ returned by LD3. Nodes are partition sets or subsets. Partition interrelations and latent confounding are abstracted away. Bidirected edge $Y\leftrightarrow\mathbf{Z}_{2}$ signifies $\mathbf{Z}_{2\notin de(Y)}$. Edges with $\cdots$ are paths of arbitrary length. Solid edges are adjacencies.
Figure 3: Baseline results for parent discovery in Sangiovese. Independence test count (Tests) is reported for constraint-based methods. Time is in seconds. Shaded regions denote 95% confidence intervals over ten replicates.
@@
Given A1, there are no descendants of $Y$ in $\mathcal{G}$ and therefore no $\mathbf{Z}_{6}$ nor $\mathbf{Z}_{2\in de(Y)}$. However, there may be $\mathbf{Z}_{2\notin de(Y)}$. Thus, $\mathcal{G}$ can contain the following seven (potentially empty) causal partitions with respect to $\{X,Y\}$: $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, $\mathbf{Z}_{4}$, $\mathbf{Z}_{5}$, $\mathbf{Z}_{7}$, and $\mathbf{Z}_{8}$ (Figure 2.A). LD3 returns the partition subsets in Figure 2.B: all $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are directly adjacent to $Y$.
Remark 2 (The observed WCDE under violations of A2).
While A2 is sufficient but not necessary for WCDE identifiability, each backdoor and frontdoor path must be blocked by at least one observable variable. If no variables on such a path are measured, then no algorithm can identify the true WCDE. Users should note that the identifiability of any direct effect measure is fundamentally limited by the observability of backdoor and frontdoor paths.
3.1 A Graphical Criterion for the Weighted CDE
Under conditions where the WCDE is identifiable by Equation 5 (Remark 1), we propose the following criterion.
Definition 8 (Graphical criterion for identifying the WCDE).
Under the causal partition taxonomy defined in Maasch et al. (2024), we define the set $\mathbf{A}_{\text{{DE}}}$ that contains all parents of the outcome:
@@ 3.1 A Graphical Criterion for the Weighted CDE
Theorem 3 ($\mathbf{A}_{\text{{DE}}}$ is a VAS for the WCDE).
$\mathbf{A}_{\text{{DE}}}$ is a valid adjustment set for the WCDE (Equation 5), satisfying the identification conditions in Definition 5.
Intuition.$\mathbf{A}_{\text{{DE}}}$ contains exactly all the parents of $Y$: confounders of $\{X,Y\}$ adjacent to $Y$ ($\mathbf{Z}_{1\in pa(Y)}$), mediators of $\{X,Y\}$ adjacent to $Y$ ($\mathbf{Z}_{3\in pa(Y)}$), and all parents of $Y$ that are marginally independent of $X$ ($\mathbf{Z}_{4\in pa(Y)}$). Thus, at least one member of every backdoor and frontdoor path is in $\mathbf{A}_{\text{{DE}}}$. This provides conditional $d$-separation of $X$ and $Y$ if and only if there is no edge $X\to Y$. Proof is in Appendix B.
Remark 3 (The role of $\mathbf{Z}_{4}$).
Note that $\mathbf{Z}_{3}$ and $Y$ can be confounded by $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, or $\mathbf{Z}_{4}$ (Figure B.1). Including $\mathbf{Z}_{4\in pa(Y)}$ in $\mathbf{A}_{\text{{DE}}}$ helps guarantee the identifiability of WCDE without requiring exact knowledge of confounding for $\mathbf{Z}_{3}$ and $Y$.
Remark 4 (Variance and Actionability).
Definition 8 defines a VAS for the WCDE in the general setting, irrespective of A1 and A2. In settings where LD3 is used to obtain $\mathbf{A}_{\text{{DE}}}$, $\mathbf{Z}$ itself constitutes a VAS (per A1 and A2). However, adjusting for all $\mathbf{Z}$ is not advised for two primary reasons:
  1. 1.
    Adjusting for all $\mathbf{Z}$ risks unnecessary adjustment, which can inflate estimator variance under finite data (Schisterman, Cole, and Platt 2009). To support statistical efficiency, we follow intuition provided by prior theorems on VAS optimality with respect to asymptotic variance, which dictate exclusion of $\mathbf{Z}_{5}$, inclusion of $\mathbf{Z}_{4}$, and generally favor control for parents of $Y$ (Rotnitzky and Smucler 2020; Henckel, Perković, and Maathuis 2022).
  2. 2.
    Ignoring causal structure limits the informativeness and actionability of fairness conclusions. Forgoing structure learning is a missed opportunity to identify potential structural mechanisms that could be redressed through interventions (e.g., policy change).
A full discussion of this topic is in Appendix C. Sections 5 and 6 provide empirical support for points (1) and (2).
@@ 4 Related Works
Discovery for CFA
Few works in causal discovery have centered on fairness objectives (Binkytė et al. 2023). These include learning Suppes-Bayes causal networks with maximum likelihood estimation (Bonchi et al. 2017) and applying PC Algorithm (Zhang, Wu, and Wu 2017). To our knowledge, this work presents the first local causal discovery method that is specifically tailored for CFA.
Local Discovery of Direct Causes
Learning the direct causes of a target has primarily garnered interest in causal feature selection (Soleymani et al. 2022). Various Markov blanket (MB) learners have been proposed, though many cannot distinguish parents from children and/or spouses (Yu et al. 2020). MB-by-MB (Wang et al. 2014), Causal Markov Blanket (CMB; Gao and Ji 2015), and Local Discovery Using Eager Collider Checks (LDECC; Gupta, Childers, and Lipton 2023) are constraint-based methods that differentiate parents and children when unambiguous over the Markov equivalence class (MEC). Like the global algorithm PC (Spirtes, Glymour, and Scheines 2001), MB-by-MB, CMB, and LDECC have exponential time complexity with respect to variable set size in the worst case. Local Discovery by Partitioning (LDP; Maasch et al. 2024) causally partitions variables around $\{X,Y\}$ using a quadratic number of CI tests with respect to variable set size; results can be post-processed to identify parents of $Y$ if A1 is imposed.
5 Empirical Validation on Synthetic Data
@@
Baselines represent a range of approaches that are available as open-source Python implementations. Local baselines are MB-by-MB (Wang et al. 2014), LDECC (Gupta, Childers, and Lipton 2023), and LDP (Maasch et al. 2024). Global baselines are PC (Spirtes, Glymour, and Scheines 2001), DirectLiNGAM (Shimizu et al. 2011), and NOTEARS (Zheng et al. 2018). Extended baseline descriptions and post-processing procedures are given in Appendix D.2. Besides LDP, all baselines assume causal sufficiency. PC, MB-by-MB, and LDECC return results in terms of the MEC. DirectLiNGAM assumes an additive noise model. DirectLiNGAM and NOTEARS assume linearity. All experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
Figure 4: Predicted parent sets, SDC, and WCDE estimates for COMPAS. Exposure is race (R; red) and outcome is general recidivism risk decile score (DS; blue). Known parents of DS are in yellow. A = age; CD = charge degree; JF = juvenile felonies; JM = juvenile misdemeanors; PC = priors count; S = sex. WCDE is reported with $p$-value ($p$) and 95% confidence intervals in brackets. All methods used $\chi^{2}$ CI tests ($\alpha=0.05$). Full results are in Figures E.3E.5.
Parent Discovery
@@
We evaluated whether LD3 can recover true parent sets using an oracle independence test on random exposure-outcome pairs in 90 unique directed acyclic Erdős-Rényi graphs (node counts in $[5...500]$) (Erdős, Rényi et al. 1960). [^6] https://r.igraph.org/ Parent F1, recall, and precision were 100%. Runtimes and total CI tests as node and edge cardinality scale are shown in Figure D.2. In Appendix D.4, we show for a linear-Gaussian SCM that WCDE estimates converged toward the true direct effect with low variance when adjusting for $\mathbf{A}_{\text{{DE}}}$ discovered with LD3 (Figures D.3, D.4).
All baselines were assessed on the Sangiovese benchmark from the bnlearn repository (Scutari 2010), a linear-Gaussian model of Tuscan grape production (Magrini, Di Blasi, and Stefanini 2017). Ten replicate datasets were sampled at $n=[250,500,1000]$. All constraint-based methods used Fisher-z tests ($\alpha=0.01$). DirectLiNGAM assumes non-Gaussian noise and was expected to underperform. LD3 was generally most performant across metrics, with LDP performing similarly (Figure 3, Table D.5). NOTEARS was significantly slower than other methods. Comparisons of LD3 to LDECC, LDP, and PC on Asia (Lauritzen and Spiegelhalter 1988) and Sachs (Sachs et al. 2005) benchmarks are in Tables D.6 and D.8. Runtime comparisons are in Figure D.2 and Tables D.7, D.9.
Estimator Variance in Finite Samples
As discussed in Remark 4, $\mathbf{Z}$ itself is a VAS under A1 and A2. However, adjusting for $\mathbf{Z}$ risks unnecessary adjustment, which can inflate the asymptotic variance of the causal effect estimator. We demonstrate the impacts of variance inflation as sample size scales in both a linear and nonlinear SCM (Figure D.5). Estimate variance using all $\mathbf{Z}$ was 7.8×7.8$\times$7.8 × to 9.6×9.6$\times$9.6 × higher than using $\mathbf{A}_{\text{{DE}}}$ in the linear SCM (Table D.2) and at least 12.6×12.6$\times$12.6 × higher in the nonlinear SCM (Table D.3).
VAS Interpretability
Structure learning can improve the interpretability of the VAS by removing irrelevant variables. In some structures, this removal can substantially reduce VAS size. On two bnlearn benchmarks with moderate to large DAGs and small $\mathbf{A}_{\text{{DE}}}$ cardinality, discovery with LD3 reduced adjustment set cardinality by at least 95.4% relative to retaining all $\mathbf{Z}$ (Table D.4).
6 Real-World Causal Fairness Analyses
We deploy LD3 for two causal fairness settings: (1) racial discrimination in criminal recidivism prediction and (2) sex-based discrimination in healthcare. We compare the results of LD3 to PC and LDECC on the basis of $\mathbf{A}_{\text{{DE}}}$ quality and computational efficiency. Causal discovery used $\chi^{2}$ CI tests and WCDE estimation used double machine learning (Chernozhukov et al. 2018). [^7] https://econml.azurewebsites.net Estimators used random forest classifiers with a 70% / 30% train-test split. [^8] https://scikit-learn.org We assumed A1 and A2. Data preprocessing is described in Appendix E. Note that all CFA results are only preliminary qualitative indicators, and further analyses should take place.
6.1 Race and COMPAS Recidivism Prediction
Background
We assessed the ability of LD3 to facilitate CFA on the COMPAS dataset from ProPublica.999github.com/propublica/compas-analysis Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) is a commercial algorithm for case management and decision support used by the US criminal justice system to assess risk of recidivism (NorthPointe 2015). ProPublica’s landmark exposé on COMPAS found that African Americans are “almost twice as likely as whites to be labeled a higher risk but not actually re-offend” (Angwin et al. 2022).
We examined racial bias in the COMPAS General Recidivism Risk model. Due to data availability, we limited our analyses to the most represented racial groups (black and white). The algorithm’s developer states that the model directly considers prior criminal history and juvenile delinquency, among other factors (NorthPointe 2015, p. 27). Our data contained multiple indicators of criminal history and juvenile delinquency, so these were used to assess the quality of causal discovery in lieu of complete ground truth. We used three significance levels for independence testing ($\alpha=0.005,0.01,0.05$) to assess stability of results (Figures 4, E.3E.5). We selected 11 features with $n=6150$ observations (2454 white, 3696 black; see Appendix E). We explored two outcomes: (1) general recidivism decile score to probe bias in the COMPAS algorithm, and (2) actual two-year recidivism, to examine factors in real outcomes.
Figure 5: Predicted parent sets, SDC, and WCDE estimates for STAR liver data. Exposure is patient sex (S; red) and outcome is receiving a liver (L; blue). Known parents of $L$ are in yellow. AE = active exception case; BT = recipient blood type; DX = diagnosis; ED = education; ET = ethnicity; EX = exception type; IA = initial age; IM = initial MELD; PM = payment method; RE = region; WE = weight. WCDE is reported with $p$-value ($p$) and 95% confidence intervals in brackets. All methods used $\chi^{2}$ CI tests ($\alpha=0.01$). Additional results are in Tables E.2E.4.
Results
At all significance levels, results from LD3, PC, and LDECC qualitatively agree that the effects of race on decile score are not fully explained by observed [^10] Note that if A2 was violated, some of the observed effects might be due to unobserved variables. This is unverifiable using ProPublica’s data alone. See Remark 2. confounding and mediation (SDC = 1) and that the WCDE is signficant ($p=0.000)$. LD3 successfully predicted that juvenile delinquency is a parent of decile score at all significance levels, while PC and LDECC never did. Age, priors count, and charge degree were nearly always predicted to be parents across methods and significance levels. For two-year recidivism, LD3 stably predicted that the WCDE of race was not significant ($p=0.87,0.68,0.77$).
In general, results for PC and LDECC were less stable than LD3. Both methods wavered between strong significance ($p=0.000$) and no significance ($p>0.4$) for the WCDE of race on two-year recidivism. LDECC parent sets were inconsistent depending on whether they were taken as the intersection or union across graphs in the MEC, requiring additional interpretation by the user. On average, PC and LDECC took 46$\times$ longer to run and performed at least 11.7$\times$ more tests than LD3 across experiments (Figure E.2).
6.2 Sex and Liver Transplant Allocation
Background